0.9(5x+15)=2.4(x+3)
Answer by stanbon(48545)

0.9(5x+15)=2.4(x+3)
Multiply both sides by 10 to get rid of the decimals.
-----
9(5x+15) = 24(x+3)
---
45x + 135 = 24x + 72
---
21x = -63
---
x = -3
=====================
Cheers,
Stan H.